lexxii:

what would be the end behavior of f(x)2x^2-5x-10/2x^2-20

AngeI:

@Hero e.e

lexxii:

\[f(x)\frac{ 2x^2-5x-10 }{ 2x^2-20}\]

lexxii:

That's better

Hero:

First factor and simplify the expression, then apply the limit as \(x \to \infty\) of \(f(x)\)

Hero:

Also apply the limit as \(x \to\) the vertical asymptote

Hero:

By \(\infty\) I am implying \(\pm \infty\)

lexxii:

this is just a practice problem. They sent a video for us to watch so we could learn and it made no sense. I just need to know what I need to do then I could figure it all out.

lexxii:

It's okay thank you Hero

Hero:

In this case, you need to find the vertical asymptote by setting the denominator equal to zero and solving for \(x\). Then apply the limit as \(x\) approaches the asymptote.

Hero:

Then you apply the limit again as \(x \to \pm \infty\) to find the horizontal asymptotes

lexxii:

by chance is the vertical asymptote \[x=\pm \sqrt{10}\] ? or did I do that completely wrong lol

Hero:

Correct

Hero:

Now find the limit as \(x \to \pm \sqrt{10}\)

lexxii:

and how do you do that? that was the farthest I actually got how to do.

Hero:

You apply the limit laws for the function. You have to apply it at least six times in this case because you have two vertical asymptotes and two horizontal asymptotes.

Hero:

So you have to find each of these: $$\lim_{x \to \sqrt{10}^+}f(x)$$$$\lim_{x \to \sqrt{10}^-}f(x)$$$$\lim_{x \to -\sqrt{10}^-}f(x)$$$$\lim_{x \to -\sqrt{10}^+}f(x)$$$$\lim_{x \to -\infty}f(x)$$$$\lim_{x \to \infty}f(x)$$

lexxii:

then after that what would I do?

Hero:

You're done after that.

lexxii:

oh okay thank you hero!

Hero:

Are you sure you know how to find them?

lexxii:

I'll be able to figure it out I have to go anyways. But thank you for the help.

Hero:

Okay, good luck